OpenStudy (thebadyboy):

CHECK PLZ A recipe calls for 10 cups of oatmeal. Benjamin wants to make 1/5 fewer servings than the recipe makes. Which expression represents how many cups of oatmeal Benjamin should use? A. 0.8 x 10 B. 1.8 x 10 C. 0.2 x 10 D. 1.2 x 10

OpenStudy (thebadyboy):

is it a or c?

OpenStudy (thebadyboy):

@jagr2713

OpenStudy (thebadyboy):

@magepker728

OpenStudy (thebadyboy):

@Mimi_x3

OpenStudy (thebadyboy):

i think c

OpenStudy (anonymous):

\[1-\frac{ 1 }{ 5 }=\frac{ 5-1 }{ 5 }=\frac{ 4 }{ 5 }\] \[\frac{ 4 }{ 5 }~of~10=?\]

OpenStudy (thebadyboy):

wat i dont get it

OpenStudy (anonymous):

one fifth fewer means four fifth

OpenStudy (thebadyboy):

i just asked if my answer is right or not

OpenStudy (anonymous):

\[\frac{ 4 }{ 5 }=0.8\]

OpenStudy (anonymous):

c is not correct.

OpenStudy (anonymous):

what about A ?

OpenStudy (thebadyboy):

yea a

OpenStudy (thebadyboy):

a is right

OpenStudy (thebadyboy):

@surjithayer